Question 7 Given vectors u = <2, 1> and v= <3,4> to compute (a) u + v (b) - (c) u-30
Answer #1

Utv 15) U-V Giren. U=<2,-17 8 V=< 3,4) : Dop (۵) U +V <2,-1) + <3,4) F2+3, -1+4) = < 5,3> = <2,-15 -(3,4) = < 2-3, -1-4> U-v